2. a) Calculate the volume under the plane z=x+2y and over the region R={(x,y)|1≤x≤2 and 3≤y≤5}.
b) Calculate the volume under the surface z=xy^2+y^3 and the region R={(x,y)|0≤x≤2 and 1≤y≤3},
1
Expert's answer
2020-11-17T16:58:42-0500

If z=f(x,y) - surface, then volume under this surface is calculated as "\int_R f(x,y)dxdy" where R is a region of xOy plane.


2.a) "\int_1^2dx\int_3^5(x+2y)dy=\int_1^2(xy+y^2)|_3^5 dx="


"=\int_1^2(5x+25-3x-9) dx=\int_1^2(2x+16) dx="


"=(x^2+16x)|_1^2=4+32-1-16=19" .


Answer: 19


2.b) "\int_0^2dx\int_1^3(xy^2+y^3)dy=\int_0^2(\frac{xy^3}{3}+\frac{y^4}{4})|_1^3dx="


"=\int_0^2(9x+\frac{81}{4}-\frac{x}{3}-\frac{1}{4})dx=\int_0^2(\frac{26x}{3}+20)dx="


"=(\frac{13x^2}{3}+20x)|_0^2=\frac{52}{3}+40=\frac{172}{3}=57\frac{1}{3}" .


Answer: "57\frac{1}{3}"